OpenStudy (anonymous):

write a system of equations with the solution (4, -3)

OpenStudy (austinl):

4x+3=19 -y+4=7

OpenStudy (austinl):

Because when you solve that out, you should get the same solutions as listed above.

OpenStudy (anonymous):

Thank you

OpenStudy (austinl):

You are welcome, it is easy enough. Try a different one on your own!

OpenStudy (anonymous):

I just needed a quick example